OpenStudy (anonymous):

Describe the solutions of the inequality in words. -10 less than or equal to 1/2c ?

OpenStudy (bahrom7893):

-10<=1/2c -20<=c c is greater than or equal to -20

OpenStudy (saifoo.khan):

\[c \ge- \frac{1}{10 \times 2}\]

OpenStudy (saifoo.khan):

Considering, \(\frac{1}{2c}\) NOT \(\frac{1}{2}c\)

OpenStudy (anonymous):

so "all real numbers less than or equal to -10" is incorrect ?

OpenStudy (bahrom7893):

what was the question asking for? please use the parenthesis.

OpenStudy (bahrom7893):

-10 less than or equal to (1/2)c or -10 less than or equal to 1/(2c) ?

OpenStudy (anonymous):

-10 less than or equal to (1/2)c and to describe the inequality in words.

OpenStudy (bahrom7893):

oh so the answer is: all real numbers greater than or equal to -20

OpenStudy (anonymous):

the answer are all rsal numbers less/equal to -20

OpenStudy (anonymous):

sorry-i mean real

OpenStudy (bahrom7893):

wait it's greater than not less than